# Implementing Evidence Acquisition:

### v2

@@ Abstract
An expert with no inherent interest in an unknown binary state can exert effort to acquire a piece of falsifiable evidence informative of it. A designer can incentivize learning using a mechanism that provides state-dependent rewards within fixed bounds. We show that eliciting a single report maximizes information acquisition if the evidence is revealing or its content predictable. This conclusion fails when the evidence is sufficiently imprecise, the failure to find it is informative, and its contents could support either state. Our findings shed light on incentive design for consultation and forecasting by showing how learning dynamics qualitatively shape effort-maximizing contracts.
@@ 1 Introduction
Often when decision makers seek informed opinions, expertise takes time to acquire and skill to interpret. For instance, senior officers deciding whether to pre-emptively respond to a potential threat or regulators deciding whether to allow an aircraft type to fly after a mishap typically rely on other specialists to learn about and communicate the socially optimal decision. [^1] In these specific applications, determine the adversary’s capabilities or investigate the incident. Given the gap between authorities making decisions and experts acquiring information about them, the former must decide how to best incentivize the latter.
This paper shows how the set of implementable expertise acquisition strategies depends jointly on the learning technology and the ability to elicit information dynamically. Our research question contrasts with similar ones from past work on information elicitation in that we ask about both dynamic learning and dynamic contracts. Our contribution is to articulate interactions between these two dimensions. We do so restricting attention to learning technologies which we refer to as finding evidence. As the name suggests, this technology reflects experts who learn until some signal arrives (i.e., the evidence is found). This class facilitates tractability, describes economically meaningful applications, and allows us to capture qualitative properties of dynamic learning technologies.
On the one hand, toward minimality, we take the agent to have no intrinsic interest in the state being learned, so that no learning would take place absent provided incentives. Outsiders, on the other hand, lack the expertise necessary to recognize the exertion of effort or interpret its outcomes. Thus, our contracting problem features both moral hazard (as the expert’s actions are unobserved) and soft information (as contracts must induce the expert to truthfully reveal what was learned). The only means to provide incentives is through rewards conditioning on the true state [^2] We note our results are maintained if rewards can depend on a signal correlated with the state, rather than the state itself. To minimize notation we suppress this possibility, but our results are equally relevant to applications where such signals are available. once revealed.
Our main results are as follows. We show that a static contract—or, more precisely, eliciting a single report from the agent—suffices to induce maximal information acquisition when (1) no information is conveyed by the failure to find evidence, (2) evidence is always conclusive, or (3) the contents of the evidence are known. But we also identify a set of parameters jointly violating these conditions for which contracts involving dynamic reports do expand the set of strategies an agent can be induced to follow.
Finding evidence learning technologies generalize the discrete-time version of the foundational Poisson bandit model. We consider an agent who chooses (privately) over time whether to exert costly effort to learn about a binary state, the initial prior over which is commonly known. Each time effort is exerted, the agent may (privately) observe an informative signal with fixed probability—reflecting the evidence having been found. Our assumption that both effort and learning are private reflects our interest in cases where evidence itself requires expertise to recognize and interpret. Many Poisson learning models from past work assume signal arrival can reveal one state (e.g., Keller et al., , 2005). In our model, while a single arrival may occur—when the sought evidence is found—we do not necessarily assume either that (a) this arrival reveals the state or (b) it can support only one state. For instance, while aviation incidents often have a single cause, this cause need not definitively imply a defect with the aircraft type. Furthermore, the conclusion the evidence would support if found might not be known in advance, either: an incident’s explanation may suggest either an underlying defect (e.g., a faulty software system) or a purely idiosyncratic issue (e.g., pilot error). Similar descriptions broadly characterize problems where experts seek some particular piece of information, aligning with our framework.
An appealing property of finding evidence settings is that the set of implementable strategies is determined by the maximum length of time that the agent can be induced to work absent evidence arrival. Indeed, the agent produces the most informative experiment precisely when working for this long. Thus, an equivalent formulation of our main question is whether varying rewards over time facilitates information production (as opposed to effort). These properties enhance tractability and intuition, making finding evidence settings a natural starting point for theoretically relating dynamic learning to incentive design.
Note that the implementation problem is trivialized if transfers or punishments can be set arbitrarily. To avoid this degeneracy and reflect practical limitations on the strength of incentives, we posit that rewards and punishments belong to some bounded range. The implementation problem may be of interest for various reasons. For one, determining the set of implementable effort strategies addresses whether a designer facing some social welfare function can induce efficiency, assuming transfers are welfare neutral (see Bergemann and Välimäki, , 2002). The implementation problem is also of interest if the information acquisition problem is of first-order importance relative to the residual value of payments or if rewards themselves are non-monetary (e.g., a promotion or special title, recognition as in Halac and Prat, (2016), a favorable recommendation, etc.).
The designer could ask the agent to report over time while acquiring information, allowing the aforementioned rewards to vary dynamically. We say a mechanism has a static implementation if it suffices to elicit just a single report of the aggregated information acquired by the agent, with the rewards provided being constant over time. Such an implementation, where rewards are provided as a function of the agent’s reported (final) belief and the true state, is known in the literature on information elicitation as a scoring rule. [^3] To our knowledge, Brier, (1950) first studied scoring rules as a way of evaluating (i.e., providing scores for) weather forecasts. In the subsequent literature, this terminology has been used to describe general reward schemes using payoffs that depend on belief reports and realized outcomes.
In one-shot interactions, the scoring rule that provides the greatest gain from exerting effort will typically depend on the prior. To see why, consider a mechanism where the agent guesses a state and is rewarded if and only if the guess is correct. If the reward when correctly guessing the initially-more-likely state is very large, then the agent might feign the discovery of evidence that confirms the prior without exerting effort. But a more modest reward in this state could get the agent to actually learn whether it is worthwhile to go against the prior. So, if the agent becomes more confident in one state, then the scoring rule that maximizes the gain from exerting effort will lower the reward provided in that state.
Applying this intuition to the dynamic setting might seem to suggest that rewards should decrease in states that the agent views as more likely the longer he exerts effort, reflecting the aforementioned comparative static for one-shot information acquisition. However, we find that this is not the case in three economically meaningful instances:
@@
    Stationary environment (Theorem 1) where beliefs absent signals are constant;
@@
    Perfect-learning environment (Theorem 2) where signals fully reveal the state;
@@
    Single-signal environment (Theorem 3) where evidence moves beliefs in one direction.
In the first case, since the agent’s belief does not change while exerting effort, the optimization problem is essentially the same at every point in time. Thus, there is no need to adjust the rewards. In the latter two cases, despite the intuition that there may be gains from “reoptimizing” rewards over time, doing so would also lower the agent’s continuation payoff from exerting effort at earlier times, weakening incentives. Essentially, reoptimizing rewards at any history redistributes the agent’s incentives to exert effort across multiple periods. No reoptimization is necessary in perfect-learning and single-signal environments.
@@
Outside of these cases, the initial intuition that more effort can be incentivized by reoptimizing the scoring rule is valid. In general, contracts with decreasing reward structures implement maximum effort (Theorem 5)—intuitively, since increasing rewards simply encourages the agent to “shirk and lie.” Perhaps more interesting in light of our other results, however, are sufficient conditions such that this type of dynamic contract strictly outperforms all scoring rules (Theorem 4). These conditions essentially boil down to a requirement that beliefs drift with a sufficiently strong violation of the second and third conditions.
@@
We emphasize that our results do not imply dynamics are irrelevant when scoring rules implement maximum effort. First, it need not be the case that the effort-maximizing scoring rule provides the strongest incentives at the prior, since incentives must be balanced as the agent’s belief changes. As we discuss, this contrasts with the static case. Second, in static settings, any level of effort that can be induced given an arbitrary mechanism can be induced by a mechanism that only requires the agent to guess a state and provides a positive reward if the guess is correct (and no reward otherwise, Li et al., (2022)). Such scoring rule formats may fail to implement an implementable amount of effort in our setting, even when the effort-maximizing contract has an implementation as a scoring rule. Specifically, as a means of encouraging the agent to exert effort at earlier times, he may be provided the option to secure a strictly positive minimum wage even if his prediction of the state is wrong.
Given the recognized importance of dynamics in information acquisition, we view our central question—i.e., how dynamic reward provision influences implementable information acquisition strategies—as a natural starting point toward developing techniques applicable to the analysis of such settings, highlighting its theoretical significance. Yet our work has notable practical implications as well—for instance, shedding light on the use of performance pay in advising contracts (e.g., with consultants, forecasters, or analysts). Often such incentives take the form of bonuses paid if the advice provided turns out to be correct (Stahl, , 2018). While idiosyncratic factors outside our simple model may determine the use and design of such contracts, understanding how to structure them to maximize the impact of incentives is still critical. Beyond our theoretical contribution, our findings provide guidelines regarding when dynamic contracts can better incentivize information acquisition. We hope these insights are practically relevant for settings where expert advice is crucial.
@@
Our paper joins a long line of work in economic theory asking how to incentivize information acquisition or experimentation. A key theoretical novelty that arises in such settings is the introduction of endogenous adverse selection since different effort choices (typically themselves subject to moral hazard) will provide the agent with different beliefs over the relevant state. This basic interaction, where an agent exerts effort under moral hazard to acquire information, has been analyzed under varying assumptions regarding the underlying information acquisition problem and contracting abilities. [^4] For instance, static information acquisition technologies where information is acquired before contracting (Crémer and Khalil, , 1992) or after (Krähmer and Strausz, , 2011), where the outcome of experimentation may be contractable (Yoder, (2022), as well as Chade and Kovrijnykh, (2016) in a repeated setting), or where decision rights but not transfers are available (Szalay, , 2005).
@@
In our model, mechanisms can condition rewards on the state as well as a report from the agent. Much of the literature on scoring rule design focuses exclusively on the elicitation of information (see, for instance McCarthy, (1956); Savage, (1971); Lambert, (2022), as well as Chambers and Lambert, (2021) for the dynamic setting). Our focus is instead on the question of how to incentivize its acquisition. To the best of our knowledge, Osband, (1989) is the earliest work to focus on this question. More recent work in economics and computer science include Häfner and Taylor, (2022); Zermeno, (2011); Carroll, (2019); Li et al., (2022); Neyman et al., (2021); Hartline et al., (2023); Whitmeyer and Zhang, (2023); Chen and Yu, (2021). Our main point of departure from this line of work stems from our focus on dynamics, and in particular the ability to write dynamic contracts. [^5] While Neyman et al., (2021); Hartline et al., (2023) and Chen and Yu, (2021) allow dynamic information acquisition, all explicitly assume contracts must be static. A related application of scoring rules is to screening forecasters, where mechanisms address initial adverse selection rather than moral hazard (see Deb et al., (2023); Dasgupta, (2023)); in particular, Deb et al., (2018) study a dynamic problem with this application, making analogous contracting assumptions to us and also describing when the solution involves a single report from the agent. [^6] A related line of work studies mechanism design with ex-post verifiability in multi-agent contexts (e.g., Deb and Mishra, , 2014; DeMarzo et al., , 2005). In this literature, Crémer, (1987) noted the potential for trivialization without restrictions in the set of mechanisms (see discussions in Deb and Mishra, (2014) and Skrzypacz, (2013)). Mylovanov and Zapechelnyuk, (2017) studies a setting with the ability to make contingent punishments, imposing restrictions on these punishments as we do.
@@
The Poisson information acquisition technology has been a workhorse for the analysis of how to structure dynamic contracts for experimentation. [^7] McClellan, (2022); Henry and Ottaviani, (2019) consider related models where information acquisition instead uses a Brownian motion technology, and an agent deciding when to stop experimenting. Bergemann and Hege, (1998, 2005) were early contributions studying a contracting problem under the assumption that a “success” reveals the state. The subsequent literature has considered variations on this basic environment (e.g., Hörner and Samuelson, (2013) relaxes commitment; Halac et al., (2016) allow for ex-ante adverse selection and transfers; Guo, (2016) considers delegation without transfers). The closest to our work is Gerardi and Maestri, (2012), who assume a Poisson arrival technology and, as in the scoring rule literature, allow for state-dependent contracts. Our information acquisition technology generalizes this technology to allow for evidence that may support either state and be inconclusive. Our results show that both modifications are necessary for dynamic mechanisms to outperform static scoring rules.
@@
On this note, Poisson bandits have been extensively utilized in strategic settings, even absent the ability to write contracts, since the influential work of Keller et al., (2005). An advantage of this setting is that it facilitates qualitative, economically-substantive properties of information acquisition and predicted behavior; a highly incomplete list of examples includes Strulovici, (2010); Che and Mierendorff, (2019); Damiano et al., (2020); Keller and Rady, (2015); Bardhi et al., (2024); Lizzeri et al., (2024). Our exercise essentially amounts to designing payoffs in a single-agent environment. Note the agent’s problem need not admit a simple stationary representation for arbitrary mechanisms in our framework. [^8] Ball and Knoepfle, (2024) study monitoring using a Poisson framework; while they allow bidirectional signals, their design problem maintains recursivity, unlike ours. This contrasts with most of the settings where payoffs are exogenous, in which case such stationarity may be crucial for tractability. Partially for this reason, our approach does not require determining the agent’s exact best response following an arbitrary dynamic contract.

### v3

# Incentivizing Forecasters to Learn:

@@ Abstract
How should forecasters be incentivized to acquire the most information when learning takes place over time? We address this question in the context of a novel dynamic mechanism design problem where a designer can incentivize learning by conditioning a reward on an event’s outcome and expert reports. Eliciting summarized advice at a terminal date maximizes information acquisition if an informative signal fully reveals the outcome or has predictable content. Otherwise, richer reporting capabilities may be required. Our findings shed light on incentive design for consultation and forecasting by illustrating how learning dynamics shape qualitative properties of effort-maximizing contracts.
@@ 1 Introduction
High-stakes decisions often rely on forecasters to assess the likelihood of future events. While sometimes forecasts take familiar forms, such as for weather or economic outlooks, others involve predictions more broadly, including national security threat assessments or medical diagnoses. This paper examines cases where the event in question is idiosyncratic, so that the only way to obtain advice is via consulting a forecaster hired to actively gather information. Furthermore, the forecaster cannot rely on passive knowledge, but must exert effort to make an informed prediction.
We start with the observation that active learning in situations such as these typically takes time. Consider, for example, a decision maker hiring an analyst to conduct due diligence before making an investment, such as acquiring a company or funding a startup. In this context, the analyst’s task is to forecast whether the investment target will meet its promised outcomes, making this determination by analyzing its complex financial records. In applications where this information is often extensive and difficult to interpret, it would be up to the analyst to decide when to stop looking for a red flag. This aspect of the information acquisition problem makes it inherently dynamic.
The problem of incentivizing forecasters to truthfully report information—rather than incentivize its acquisition—has been extensively studied in the literature on scoring rule design, starting with Brier, (1950) which considered how to evaluate (i.e., provide scores for) weather forecasts. Building on this line of work, our paper addresses the design of incentives to best encourage both information acquisition and truthful revelation, rather than just the latter. In particular, the literature on forecasting often emphasizes that a primary motivator for forecasters is reputational concerns (Marinovic et al., , 2013). If the forecaster is employed by a decision maker, the decision maker’s recommendation has a fixed value to the forecaster because it can influence perceptions. This allows the principal to incentivize learning by conditioning the endorsement on the expert’s prediction and the realized outcome. [^1] See Petropoulos et al., (2022) for a recent survey on forecasting, including a discussion of the relevance of scoring rules for forecaster incentives. Gneiting and Raftery, (2007) provides theoretical background on scoring rules, while Frongillo and Waggoner, (2023) provides a more recent survey. Following this line of work, we consider a designer able to provide a reward of fixed value (i.e., a score) as a means of incentivizing the forecaster.
Our contribution is to identify learning technologies for which a single prediction, made after all information is acquired, not only induces truthful reporting but also maximizes incentives for information acquisition. These results clarify when richer contracting environments are more conducive to informed forecasts. In particular, some relationships—such as when an outside consultant is hired at arm’s length and restricted to only provide a single report at some prespecified date—may limit the scope for reports to be provided at any time. Closer relationship may make constant reporting feasible, but could introduce other costs. Whether such costs are worthwhile depends on whether dynamic communication is necessary to maximize information acquisition incentives.
Our goal is to provide a framework sufficiently general to offer broad, qualitative insights related to whether scoring rules maximize the incentives for information acquisition. Toward that end, we utilize a tractable framework enabling sharp comparisons despite known complexities due to the combination of dynamic information acquisition and a rich contracting space. [^2] To briefly comment on some of the potential complexities, we mention that the agent’s decision problem need not be stationary for general contract choices of the designer. By contrast, much work on costly information acquisition relies heavily on such stationarity to deliver tractability, requiring us to develop new techniques in order to speak to contract design. Specifically, we study a version of a discretized Poisson bandit learning technology. We consider an agent who chooses (privately) over time whether to exert costly effort to learn about an uncertain future binary event. We refer to the outcome of this event as the state, and assume the initial prior over it is commonly known.
When exerting effort, the agent may (privately) observe two kinds of signals, each (in our main model) with fixed probability: (a) a “null signal” or (b) a “Poisson signal.” We interpret a Poisson signal arrival as the event that the forecaster has observed some particular sought-after information. This would correspond to a red flag in the due diligence application described above. In a national security application, this would correspond to a determination of an adversary’s capabilities. We emphasize that a Poisson signal need not reveal the outcome or suggest only one outcome. For instance, a red flag might identify which outcomes might lead the investment to fail, but not definitively reveal if any of these outcomes will arise. In a national security context, a forecaster might determine whether an adversary has the capabilities to launch an attack on some specified date, but not definitively whether they intend to do so. Similar descriptions broadly characterize problems where experts seek some particular piece of information, aligning with our framework.
Our assumption that both effort and learning are private reflects our interest in cases where the sought-after signal itself requires expertise to recognize and interpret. To make the comparison as clear as possible, we aim to keep the contracting environment with a single elicitation at a prespecified date otherwise identical to the general case except for the ability to receive reports from the agent at any time. Given this, the designer’s problem in either contracting environment is to decide how to provide the endorsement or recommendation as a function of (a) the forecaster’s prediction and (b) the outcome that arises.
We mention that the design of contracts to elicit advice in dynamic settings has a much shorter tradition. An important paper sharing this focus with us is Deb et al., (2018), studying the question of how to design such dynamic contracts when aiming to screen forecaster ability without moral hazard considerations. While we share their assumptions on the principal’s contracting ability, the most substantive difference between our exercises is that we consider dynamic moral hazard instead of initial adverse selection.
In the aforementioned literature on information elicitation, a scoring rule is a contract where the recommendation is provided as a function of (a) a single forecaster belief report, and (b) the realized outcome. Our main results identify three cases under which a scoring rule can implement maximal effort in the dynamic setting:
@@
    If no information is conveyed by the failure to find a red (or green) flag: That is, a Stationary environment (Theorem 1) where beliefs absent signals are constant;
@@
    If a red (or green) flag reveals the future outcome perfectly: That is, a Perfect-learning environment (Theorem 2) where signals fully reveal the state;
@@
    If only a red flag can arrive: That is, a Single-signal environment (Theorem 3) where a Poisson signal arrival always moves beliefs in one direction.
Outside of these cases, dynamic reports may expand the set of strategies an agent can be induced to follow. In particular, we identify a set of parameters jointly violating these conditions for which richer contracts are necessary.
We briefly describe why one might expect dynamic contracts to be necessary in our problem. We observe that in one-shot interactions, the scoring rule that provides the greatest gain from effort will typically depend on the prior. To see why, consider a mechanism where the agent guesses a state and is rewarded if and only if the guess is correct. If the reward when correctly guessing the initially-more-likely state is very large, then the agent might feign the arrival of a Poisson signal that confirms the prior without exerting effort. But a more modest reward in this state could get the agent to actually learn whether it is worthwhile to go against the prior. So, if the agent becomes more confident in one state, then the scoring rule that maximizes the gain from exerting effort will lower the reward provided in that state. Now, it is immediate that such adjustments are unnecessary for the stationary environment, since beliefs do not change prior to a Poisson signal arrival. Thus, the optimization problem is essentially the same at every point in time, and it is straightforward to show there is no need to adjust rewards over time.
However, for the other environments (perfect learning and single-signal), taking this intuition to the dynamic setting might seem to suggest that rewards should decrease in states that the agent views as more likely the longer he exerts effort. Indeed, we discuss in Section 3.2 how similar intuition does imply rewards should be lower in initially more likely states when information acquisition is one-shot. But despite intuition that there may be gains from adjusting rewards over time, doing so in the dynamic case lowers the agent’s continuation payoff from exerting effort at different times, weakening incentives. Adjusting rewards at any history to maximize the incentives to exert effort at the agent’s new belief redistributes the agent’s incentives to exert effort across multiple periods. This observation drives the result that no adjustments is necessary in the perfect-learning and single-signal environments. Intuitively, contracts varying rewards over time necessarily entail weaker incentives for some potential expert beliefs in order to maintain incentive compatibility. For these environments, any contract adjusting rewards over time can be replaced by a (static) scoring rule providing weakly stronger incentives at all times.
@@
In general, effort-maximizing contracts can be determined numerically via linear programming. We describe this in Section 3.1. We caution, however, that while the numerical procedure does illustrate the structure of effort-maximizing contracts—both for summarized and unrestricted advice—it does not necessarily clarify when the unrestricted solution can be implemented as a scoring rule. In other words, while reassuring that one can compute effort-maximizing contracts, these results provide limited insights into their qualitative properties. We do show that contracts with decreasing reward structures implement maximum effort (Theorem 5)—intuitively, since increasing rewards simply encourages the agent to “shirk and lie.” More directly relevant to our main message, however, are sufficient conditions such that this type of dynamic contract strictly outperforms all scoring rules (Theorem 4). These conditions essentially require that beliefs drift with a sufficiently strong violation of the second and third conditions. This result shows that the ability for scoring rules to maximize effort should not be taken for granted, in general.
@@
Our results do not imply dynamics are irrelevant when scoring rules implement maximum effort. First, it need not be the case that the effort-maximizing scoring rule provides the strongest incentives at the prior, since incentives must be balanced as the agent’s belief changes, as alluded to above. Second, in static settings, any level of effort that can be induced given an arbitrary mechanism can be induced by a mechanism that only requires the agent to guess a state and provides a positive reward if the guess is correct (and no reward otherwise, Li et al., (2022)). Such scoring rule formats may fail to implement an implementable amount of effort in our setting, even when the effort-maximizing contract has an implementation as a scoring rule. Specifically, as a means of encouraging the agent to exert effort at earlier times, he may be provided with an intermediate option to secure a strictly positive minimum wage even if his prediction of the state is wrong. This observation highlights a sharp contrast with the findings in the static model by Szalay, (2005), which provides intuition for why optimal contracts should only offer extreme options to make the agent sufficiently invested in information acquisition. The differing conclusion is due to the fact that beliefs change over time in our setting—hence, an option that is “intermediate” early on may be relatively “extreme” later. Even when static scoring rules suffice, effort-maximization requires contracts to anticipate the evolution of beliefs.
@@
Our paper joins a long line of work in economic theory asking how to incentivize information acquisition or experimentation. A key theoretical novelty that arises in such settings is the introduction of endogenous adverse selection since different effort choices (typically themselves subject to moral hazard) will provide the agent with different beliefs over the relevant state. This basic interaction, where an agent exerts effort under moral hazard to acquire information, has been analyzed under varying assumptions regarding the underlying information acquisition problem and contracting abilities. [^3] For instance, static information acquisition technologies where information is acquired after contracting (Krähmer and Strausz, , 2011) or where the outcome of experimentation may be contractable (Yoder, (2022), as well as Chade and Kovrijnykh, (2016) in a repeated setting).
@@
As alluded to above, much of the literature on scoring rule design focuses exclusively on the elicitation of information (see, for instance McCarthy, (1956); Savage, (1971); Lambert, (2022), as well as Chambers and Lambert, (2021) for the dynamic setting). To the best of our knowledge, Osband, (1989) is the earliest work sharing our focus on the question of incentivizing the acquisition of information. Other work relevant to the application of forecasting is Elliott and Timmermann, (2016), which reviews statistical properties of forecasting models. Aside from Deb et al., (2018), other papers focused on the problem of screening forecasters include Deb et al., (2023); Dasgupta, (2023). More recent work in economics and computer science related to scoring rules include Häfner and Taylor, (2022); Zermeno, (2011); Carroll, (2019); Li et al., (2022); Neyman et al., (2021); Hartline et al., (2023); Whitmeyer and Zhang, (2023); Chen and Yu, (2021); Bloedel and Segal, (2024). Our main point of departure from this line of work stems from our focus on dynamics. [^4] While Neyman et al., (2021); Hartline et al., (2023) and Chen and Yu, (2021) allow dynamic information acquisition, all explicitly assume contracts must be static. Bloedel and Segal, (2024) study dynamic contracts, but their framework models dynamics across different agents, whereas our model focuses on dynamics in which a single agent acquires information over time.
@@
The Poisson information acquisition technology has been a workhorse for the analysis of how to structure dynamic contracts for experimentation. [^5] McClellan, (2022); Henry and Ottaviani, (2019) consider related models where information acquisition instead uses a Brownian motion technology, and an agent deciding when to stop experimenting. Bergemann and Hege, (1998, 2005) were early contributions studying a contracting problem under the assumption that a “success” reveals the state. The subsequent literature has considered variations on this basic environment (e.g., Hörner and Samuelson, (2013) relaxes commitment; Halac et al., (2016) allow for ex-ante adverse selection and transfers; Guo, (2016) considers delegation without transfers). The closest to our work is Gerardi and Maestri, (2012), who assume a Poisson arrival technology and, as in the scoring rule literature, allow for state-dependent contracts. Our information acquisition technology generalizes this technology to allow for Poisson signals that may support either state and be inconclusive. Our results show that both modifications are necessary for dynamic mechanisms to outperform static scoring rules.
@@
On this note, Poisson bandits have been extensively utilized in economic settings since the influential work of Keller et al., (2005). An advantage of this framework is that it facilitates qualitative, economically-substantive properties of information acquisition and predicted behavior; a highly incomplete list of examples includes Strulovici, (2010); Che and Mierendorff, (2019); Damiano et al., (2020); Keller and Rady, (2015); Bardhi et al., (2024); Lizzeri et al., (2024). Our exercise essentially amounts to designing a (possibly dynamic) single-agent decision problem. Note the agent’s problem need not admit a simple stationary representation for arbitrary mechanisms in our framework. [^6] Ball and Knoepfle, (2024) study monitoring using a Poisson framework; while they allow bidirectional signals, their design problem maintains recursivity, unlike ours. This contrasts with most of the settings where payoffs are exogenous, in which case such stationarity may be crucial for tractability. Partially for this reason, our approach does not require determining the agent’s exact best response following an arbitrary dynamic contract.